OpenStudy (anonymous):

Which line is a graph of the equation: –2x + 5y = 10?

OpenStudy (anonymous):

I WILL GIVE MEDALS!!

pooja195 (pooja195):

Hey there! Welcome to openstudy! Are we supposed to find the intercepts? Or are we supposed to match to a graph?

pooja195 (pooja195):

The question doesn't make much sense

OpenStudy (texaschic101):

I think there is supposed to be a graph

OpenStudy (anonymous):

Oh I forgot hold on.

OpenStudy (anonymous):

How do I post it??

OpenStudy (texaschic101):

to find y intercept, sub in 0 for x and solve -2x + 5y = 10 -2(0) + 5y = 10 5y = 10 y = 2.....so your y intercept is (0,2) to find x intercept, sub in 0 for y -2x + 5y = 10 -2x + 5(0) = 10 -2x = 10 x = -5....so your x intercept is (-5,0) now what graph has the intercepts of (0,2) and (-5,0) ??

OpenStudy (anonymous):

I was thinking line D.

OpenStudy (texaschic101):

no...not D. which line crosses the x axis at (-5,0) ? which line crosses the y axis at (0,2) ??

OpenStudy (anonymous):

A

OpenStudy (texaschic101):

yep ^^ you got it :)

OpenStudy (anonymous):

@texaschic101 Sorry I blew up at you earlier.

OpenStudy (texaschic101):

its understandable....too much homework can do that to ya...lol

OpenStudy (anonymous):

Are we cool??

OpenStudy (texaschic101):

of course...if I had $1 for everybody I " blew up " at...I would be rich

OpenStudy (anonymous):

Cool. Have a great night!! Bye.